# Fast and Accurate Collimator-Detector Response Compensation in High-Energy SPECT Imaging with 1D Convolutions and Rotations

### v1

@@ 2.1.2 Image Reconstruction
All acquired image data in subsequent sections are reconstructed using PyTomography with the maximum likelihood expectation maximization (MLEM) algorithm; this was chosen over the ordered subset expectation maximization algorithm (OSEM) because of the low count statistics for each projection. System matrices employed attenuation correction using attenuation maps derived from acquired computed tomography (CT) images. As described by Zekun et al. [38], SPECT with $\alpha$-RPTs often have significant stray radiation-related noise component due to the proportionally low count rate. The imaging system equation is thus given by
$\bar{y}=Hx+\bar{\Psi}+\hat{s}$ (20)
where $\bar{y}$ is the expectation of the acquired count data $y$, $H$ is the system matrix, $x$ is the estimated 3D count distribution, $\bar{\Psi}$ is the expectation of the stray radiation-related noise, and $\hat{s}$ is the scatter estimate. $\bar{\Psi}$ is assumed to be the same for all bins in a given energy window and proportional to the acquisition time, and can be measured by taking a blank scan using the same energy windows as the patient acquisition, averaging the number of counts in each valid bin, and scaling by the acquisition time used in the clinical protocol. Since the dual energy window (DEW) technique is used for scatter estimation, the stray radiation in that window also must be accounted for: the scatter estimation is thus given by
$\hat{s}=\left(\frac{w_{p}}{w_{l}}\right)\left(S_{G}[g_{l}]-\bar{\Psi_{l}}\right)$ (21)
where $w_{p}$ is the width of the primary window, $w_{l}$ is the width of the lower window, $S_{G}$ is a Gaussian smoothing kernel, $g_{l}$ is the acquired count data in the lower energy window, and $\Psi_{l}$ is the expectation of the mean stray-radiation noise in the lower energy window. The smoothing kernel is used here due to the high noise present. One of the dangers of Equation 21 is that it is capable of yielding negative values in Equation 20; in practice this rarely occurred, and all negative values were set to zero.
When used in image reconstruction, the size of the 2D PSF kernels was always set to $N-1\times N-1$ where $N$ is the number of voxels along the largest direction in the reconstructed image. For the 1D-R model, the kernels corresponding to $K_{G}$ and $K_{B}$ were of size $N-1$, while the kernel corresponding to $K_{t}$ was of size $\lceil\sqrt{2}N\rceil+n$ to account for the diagonal, where $n$ is of integer value 0 or 1 to make the kernel size odd. It should be noted that for PSFs much larger than the dimensions of the image, the size of the kernels would need to cover a $12N-1\times 2N-$ 1 area to account for contributions from voxels on one edge of the image to detector elements on the opposite edge; in practice, the PSFs were not that large and the dimensions chosen were sufficient.
@@ 2.2 Validation and Timing
A 440 keV point source was simulated using the SIMIND Monte Carlo program [37] at 1100 positions that linearly varied from 0 cm to 58.44 cm. The detector pixel size was $0.24~{}\mathrm{cm}\times 0.24~{}\mathrm{cm}$ with $255\times 255$ pixels. The simulated collimator corresponded to a Siemens high energy configuration with a hexagonal shape, a hole width of 0.40.40.4~{}0.4cm and a hole length of 5.975.975.97~{}5.97cm. The intrinsic resolution of the detector was also included in the simulation and was assumed to be 0.380.380.38~{}0.38cm at 140 keV, representative of a Symbia system with $3/8$” crystal length. The PSFs at all 1100 distances were used to generate $K_{\mathrm{true}}^{x,y;d}$ in Equation 19 during reconstruction by selecting the PSF closest to the actual source-detector distance.
Twelve of these PSFs at source detector distances of 111~{}1cm and every 555~{}5cm from 555~{}5cm to 555555~{}55cm were used as data to fit the 1D-R model. Fitting was performed in three steps: (i) the Gaussian parameters were optimized using the Adaptive Moment Estimation (ADAM) algorithm for $1\cdot 10^{4}$ iterations and a learning rate of $10^{-2}$, (ii) all the parameters were simultaneously optimized using ADAM for $1.5\cdot 10^{4}$ iterations and a learning rate of $10^{-3}$.
The computational time for the 1D-R model (Equation 8) and 2D model (Equation 19) were bench-marked on CPU and GPU for by applying the operators to matrices of four different sizes: $64^{3}$, $128^{3}$, $196^{3}$, and $256^{3}$. Each matrix of size $N^{3}$ was filled with random uniform numbers between 0 and 1. Each experiment used conditions similar to SPECT imaging, where the $d$ axis varied from 000~{}cm to 505050~{}50cm, and the $x$ and $y$ axes varied from $-30.74~{}$cm to 30.7430.7430.74~{}30.74cm. Use of standard convolution and fast fourier transform (FFT) based convolution were compared for each method.
@@ 2.3 SPECT Studies
2.3.1 Phantom Study
Application of (i) standard Gaussian (ii) 1D-R (iii) 2D PSF modeling was evaluated for GPU reconstruction of acquired 225Ac data. A cylindrical phantom with spheres of diameter 606060~{}60mm, 282828~{}28mm, and 222222~{}22mm with an initial sphere activity concentration of 1.37 kBq/mL was filled at a 10:1 source to background ratio. 34 SPECT acquisitions of the phantom were taken in sequence on a Symbia T2 SPECT/CT system (Siemens Healthineers, USA) with the following settings: $128\times 128$ pixels at $4.82~{}\mathrm{mm}\times 4.82~{}\mathrm{mm}$ resolution, 96 projection angles, high energy collimators, and 60 s acquisition time per projection. For reconstruction of the 440440440~{}440keV emission from the 213Bi daughter, two energy windows were configured for acquisition: (i) photopeak window centered at 440 keV with a width of 20% and (ii) a lower scatter window used for DEW scatter correction centered at 374374374~{}374keV with a width of 12.5%. A blank scan with identical parameters was acquired to obtain the mean-stray radiation noise in each energy window. Two noise levels of the data were considered:
@@ 2.3.1 Phantom Study
    One of the 34 scans was used for data reconstruction. The number of acquired counts in this scenario corresponds to an expected clinical 225Ac scan of a patient undergoing 225Ac-PSMA-617 therapy with injection of 888~{}8MBq [23] and is scanned for 2.5 min per projection at a time point between 0-72 hr post injection.
@@
    The counts from all 34 scans were summed together and used for image reconstruction. In this case, no scatter blurring was used in Equation 21 since the noise level was low. Because the detectors had an approximately equal radial path around the phantom for each scan, this corresponds to a single scan with a high count rate.
Although 96 projection angles were acquired, only 32 were used since this was similar to the corresponding 225Ac patient acquisition in the next section. All images were reconstructed with MLEM for up to 100 iterations.
@@ 2.3.2 Patient Study
A patient receiving 225Ac-PSMA-617 therapy with an injected activity of 8 MBq was imaged 20.5 hours post injection (first cycle). SPECT data were acquired on a Discovery 670 Pro SPECT/CT (GE Healthcare, USA) with a high energy general purpose collimator. 30 projections (15 per head) were acquired for 150 s using an identical photopeak window to the phantom acquisition, and 358.15 keV - 395.85 keV lower window. Since this GE scanner has a different CDR than the Siemens scanner used in prior examples, PSF data were regenerated in SIMIND using a high energy general purpose GE collimator and another 1D-R model was fit as before. In GE scanners, the hexagonal bores in the collimator are rotated by $90^{\circ}$ compared to the Siemens collimator, so the angles in $\Theta_{t}$ were adjusted to compensate for this. The acquired data were reconstructed using (i) standard Gaussian (ii) 2D and (iii) 1D-R PSF modeling using MLEM for up to 100 iterations. Reconstructed images were filtered using a Gaussian filter with a full width half max (FWHM) of 333~{}3cm. Two high uptake bone lesions were segmented on the 100th iteration reconstruction using 3D Slicer [39]. Three lesion ROIs were segmented by a physician on a pre-therapeutical PET image using the PET Edge+ tool of MIM v7.2.1 (MIM Software Inc., USA). A spherical ROI of diameter 101010~{}10cm was placed in a low uptake region in the center of the patient so that the contrast, defined as the mean uptake ratio between the lesion ROIs and background ROI, could be obtained. The mean number of counts and contrast for each lesion ROI were then evaluated for each iteration of MLEM; before the statistics were computed, the 333~{}3cm FWHM Gaussian filter was applied to the image.
@@ 3 Results
Figure 2 shows the Monte Carlo simulated PSF data compared to the fit obtained via Equation 8. The curve fit is a reasonable approximation to the PSF at most distances, but is unable to capture features at larger distances, such as the intensity of the tails at 50 cm. The model also consistently under predicts the intensity at the center of the PSF, and this effect is more prevalent at large distances.
@@
Figure 2: Comparison between 225Ac Monte Carlo simulated PSF data and the corresponding fit from the 1D-R model at different source-detector distances. From top to bottom, (i) 2D profiles of Monte Carlo PSF data, (ii) 2D profiles of 1D-R model, (iii) central vertical profile of Monte Carlo and fitted PSFs, (iv) central horizontal profile of Monte Carlo and fitted PSFs; the Monte Carlo data is shown in black scatter points, while the fitted data is shown as a solid red line.
@@
Timing benchmarks for PSF modeling using the 1D-R, 1D-R (FFT), 2D, and 2D (FFT) are shown in Figure 3. CPU implementation yields no benefits with the proposed 1D-R model, and performs fastest with FFT-based 2D convolutions and slowest using regular 2D convolutions. The GPU implementation is faster than the CPU implementation for all methods, and yields computational benefits when the proposed 1D-R model is used. With a matrix size of $128^{3}$, use of the 1D-R method is over three times faster than 2D (FFT) PSF modeling. For small matrix sizes ($64^{3}$) there is no computational speed-up using the 1D-R method, and for large matrix sizes ($256^{3}$) the relative time difference between the 1D-R and 2D (FFT) begins to decrease, approaching only a two times speed advantage.
@@
Figure 3: CPU and GPU required run time of the 1D-R and 2D PSF models for different matrix sizes. Each implementation is evaluated using standard and FFT based convolution. The time shown on top of each bar in units of ms.
@@
PyTomography reconstructed 225Ac phantom images are shown in Figure 4. The time required for 100 iterations of MLEM for the low count data was 110110110~{}110s (Gaussian PSF), 146146146~{}146s (1D-R PSF) and 377377377~{}377s (2D PSF). Reconstructions using the 1D-R and 2D methods are almost indistinguishable qualitatively. All 34 low count acquisitions were then reconstructed and the 1D-R and 2D methods were quantitatively compared; the differences between the 2D and 1D-R were $(1.25\pm 1.25)\%$, $(0.60\pm 0.69)\%$, and $(1.90\pm 0.18)\%$ from smallest to largest sphere respectively. For comparison, the variability in counts between separate noise realizations for the 2D model were $54.4\%$, $33.0\%$, and $9.8\%$ from smallest to largest sphere, respectively.
@@
Figure 4: Axial slices from reconstruction of low-count (top row) and high-count (bottom row) 225Ac phantom data acquired on a Symbia T2 SPECT system. Reconstructed SPECT and CT images are superimposed; the boundaries of each sphere can be seen on the CT. Images were reconstructed with the reference 2D, Gaussian, and proposed 1D-R PSF modeling using the acquired Monte Carlo data. The displayed low count images were post-smoothed using a 3D Gaussian function with a 3 cm FWHM. The plots on the right show the average counts per voxel in the small (sml) medium (med) and large (lrg) spheres as a function of iteration number; the 2D is shown in black, the Gaussian is shown in blue, and the 1D-R is shown in red.
PyTomography reconstructed 225Ac-PSMA-617 patient images are shown in Figure 5. The time required for 100 iterations of MLEM was $38.5$ s (Gaussian PSF), 165.6165.6165.6~{}165.6s (1D-R PSF) and 351.5351.5351.5~{}351.5s (2D PSF). Qualitatively, the Gaussian PSF yields less counts in the uptake ROIs, and also greater background counts, for example, in the vertebrae. The count differences between the 1D-R and 2D methods after 100 iterations were 1.271.271.27~{}1.27(lesion 1), 0.480.480.48~{}0.48(lesion 2) and $2.03~{}\%$ (lesion 3). These differences are comparable to those observed in the phantom study, and were significantly less than the differences between the Gaussian and 2D methods of $-42.3~{}\%$ (lesion 1), $-29.1~{}\%$ (lesion 2) and $-44.3~{}\%$ (lesion 3). The differences in lesion-background contrast between the 1D-R and 2D methods after 100 iterations were $-0.18~{}\%$ (lesion 1), $-0.96~{}\%$ (lesion 2) and $0.56~{}\%$ (lesion 3). These were smaller in magnitude than the contrast differences between the Gaussian and 2D methods of $-51.3\%$ (lesion 1) and $-23.3~{}\%$ (lesion 2) and $-56.7~{}\%$ (lesion 3). Based on the line plots in Figure 5, the Gaussian PSF yields convergence with less iterations; this observation is similar to how an absence CDR modeling leads to faster convergence in traditional low and medium energy SPECT. In general, the smaller the spatial extent of the PSF, the faster convergence is attained in MLEM.
@@
Figure 5: Coronal slices of 68Ga-PSMA-617 (top row, left) and 225Ac-PSMA-617 patient data (top row, right), counts in bone lesion ROIs as a function of iteration number (middle row) and signal to background ratios in the bone lesion ROIs (bottom row). The Background ROI was a 10 cm sphere placed in the center of the patient with little activity. The 68Ga-PSMA-617 image was reconstructed using QClear on the vendor’s software. 225Ac images were reconstructed in PyTomography using (i) reference 2D, (ii) Gaussian, and (iii) the proposed 1D-R model with 100 iterations of MLEM. The displayed reconstructed SPECT images were post filtered using a 3D 3 cm FWHM Gaussian. For computation of the data in the line plots, the image at each iteration was post-filtered using a 3 cm FWHM 3D Gaussian filter before the statistic was computed. Arrows pointing to the three bone lesions are shown in red (lesion 2 is directly on top of its annotation). There was no uptake in the right kidney due to organ failure.

### v2

@@ 2.1.2 Image Reconstruction
All acquired image data in subsequent sections are reconstructed using PyTomography with the ordered subset expectation maximization algorithm (OSEM); when one subset is used, the algorithm is denoted as maximum likelihood expectation maximization (MLEM). System matrices employed attenuation correction using attenuation maps derived from either (i) ground truth attenuation maps from MC simulations or (ii) acquired computed tomography (CT) images in physical studies. The imaging system equation is given by
$\bar{y}=Hx+\hat{s}$ (20)
where $\bar{y}$ is the expectation of the acquired count data $y$, $H$ is the system matrix, and $x$ is the estimated 3D count distribution. $\hat{s}$ is an additive term that accounts for phantom or patient scatter, events that back-scattered in the detector from higher energy emission peaks, and stray background radiation; the presence of stray background radiation for low count SPECT has been described by Zekun et al. [40]. For all cases in this paper, the triple energy window (TEW) technique is used to estimate $\hat{s}$.
When used in image reconstruction, the size of the 2D PSF kernels in the 2D-MC method was always set to $N-1\times N-1$ where $N$ is the number of voxels along the largest direction in the reconstructed image. For the 1DR-MC method, the kernels corresponding to $K_{G}$ and $K_{B}$ were of size $N-1$, while the kernel corresponding to $K_{t}$ was of size $\lceil\sqrt{2}N\rceil+n$ to account for the diagonal, where $n$ is of integer value 0 or 1 to make the kernel size odd. It should be noted that for PSFs much larger than the dimensions of the image, the size of the kernels would need to cover a $12N-1\times 2N-$ 1 area to account for contributions from voxels on one corner of the image to detector elements on the opposite corner; in practice, the PSFs were not that large and the dimensions chosen were sufficient.
2.1.3 Image Quality Metrics
To compare images reconstructed using the three different PSF models, two image metrics are used. The first metric is recovery coefficients (RCs), defined as
$\mathrm{RC}\equiv\frac{A_{\mathrm{pred}}}{A_{\mathrm{true}}}$ (21)
where $A_{\mathrm{pred}}$ corresponds to the predicted activity concentration in a region of interest (ROI), and $A_{\mathrm{true}}$ corresponds to the true activity in that region. In MC simulations consisting of spheres with infinite resolution, the ROIs on the reconstructed voxel grid are segmented such that voxels inside ROI are assigned 1 while voxels outside are assigned 0. The calibration factors necessary to convert counts to activity units are obtained via point source simulations with the same detector parameters. For the physical data, where calibration factors are not available, average counts are instead reported. The second metric used is contrast, defined as
$\mathrm{Contrast}\equiv\frac{C_{\mathrm{hot}}}{C_{\mathrm{bkg}}}$ (22)
where $C_{\mathrm{hot}}$ is the number of counts in a hot ROI and $C_{\mathrm{bkg}}$ is the number of counts in the background.
@@ 2.2 Validation and Timing
Point sources with energies of 208 keV, 364 keV, and 440 keV (representative of the typical primary energies acquired in 177Lu, 131I, and 225Ac imaging respectively) were simulated using the SIMIND Monte Carlo program [37] at 1100 positions that linearly varied from 0 cm to 58.44 cm. The detector pixel size was $0.24~{}\mathrm{cm}\times 0.24~{}\mathrm{cm}$ with $255\times 255$ pixels. For 177Lu, the simulated collimator corresponded to a Siemens low energy high resolution (LEHR) configuration. While typical clinical protocols use a medium energy (ME) configuration to minimize septal penetration and scatter, the choice of collimator here is merely to demonstrate the feasibility of using a LEHR collimator with an advanced PSF model in reconstruction. A Siemens high energy (HE) configuration was used for 131I and 225Ac, which is standard for clinical protocol. The intrinsic resolution of the detector was also included in the simulation and was assumed to be 0.380.380.38~{}0.38cm at 140 keV, representative of a Symbia system with $3/8$” crystal length. The collimator cover and backscatter components of the detector were modeled via 0.1 cm of aluminum and 6.6 cm of pyrex respectively; inclusion of these parameters is necessary to capture further scattering of detected photons present in real systems.
Twelve of the simulated PSFs per isotope/collimator configuration (at source-detector distances of 111~{}1cm and every 555~{}5cm from 555~{}5cm to 555555~{}55cm) were used as data to fit the 1DR-MC model. Fitting was performed in three steps: (i) the Gaussian parameters were optimized using the Adaptive Moment Estimation (ADAM) algorithm for $1\cdot 10^{4}$ iterations and a learning rate of $10^{-2}$, (ii) all the parameters were simultaneously optimized using ADAM for $1.5\cdot 10^{4}$ iterations and a learning rate of $10^{-3}$. The PSFs at all distances were used to build the 2D-MC models.
The computational time for the 1DR-MC and 2D-MC models were bench-marked on CPU and GPU for by applying the 225Ac operators to matrices of four different sizes: $64^{3}$, $128^{3}$, $196^{3}$, and $256^{3}$. Each matrix of size $N^{3}$ was filled with random uniform numbers between 0 and 1. Each experiment used conditions similar to SPECT imaging, where the $d$ axis varied from 000~{}cm to 505050~{}50cm, and the $x$ and $y$ axes varied from $-30.74~{}$cm to 30.7430.7430.74~{}30.74cm. Use of standard convolution and fast fourier transform (FFT) based convolution were compared for each method.
@@ 2.3 SPECT Studies
2.3.1 Monte Carlo Phantom Studies
To demonstrate the versatility of the proposed method, application of the 2DG, 2D-MC, and 1DR-MC models were evaluated for reconstruction of MC simulated 177Lu, 131I, and 225Ac projection data. Each case consisted a cylindrical phantom with spheres of diameter 606060~{}60mm, 282828~{}28mm, and 222222~{}22mm, and a source to background ratio of 10:1. The activity concentrations in the spheres of 200200200~{}200Bq/mL (225Ac) and 111~{}1MBq/mL (177Lu and 131I) were selected based on the expected activity concentrations in patients. The SIMIND detector and collimator parameters were identical to the point source simulations, but data were now collected at a matrix size of $128\times 128$ pixels with $4.8~{}\mathrm{mm}\times 4.8~{}\mathrm{mm}$ resolution. 96 projection angles at 15 s/projection were collected for 177Lu and 131I, while 32 projection angles at 45000 s/projection were collected for 225Ac; while the 177Lu and 131I configurations were representative of a typical clinical acquisition, the 225Ac configuration was such that the number acquired counts was significantly higher. The energy windows used for TEW scatter correction of each isotope are shown in Table 1. Projection data were reconstructed with OSEM for up to 20 iterations with 8 subsets (177Lu and 131I) and up to 200 iterations with 1 subset (225Ac). Calibration factors were used to convert images to units of MBq/mL, and were obtained via corresponding point source simulations with the same detector configuration. The background region used to compute contrast corresponded to four 30 mm spherical ROIs placed in the warm region.
Table 1: Energy windows employed for all image acquisitions and reconstructions in this paper to the nearest decimal in keV.
  • Isotope Photopeak Lower Upper
    177Lu (MC Phantom) 187.2-228.8 169.4-187.2 228.8-252.9
    131I (MC Phantom) 332.0-405.0 310.0-332.0 405.0-427.0
    225Ac (MC Phantom) 396.0-484.0 374.0-396.0 484.0-506.0
    225Ac (Physical Phantom) 396.0-484.0 352.0-396.0 484.0-572.0
    225Ac (Patient) 396.0-484.0 358.2-395.9 484.0-506.0
2.3.2 Physical Phantom Studies
Application of the 2DG, 2D-MC, and 1DR-MC models were evaluated for reconstruction of acquired 225Ac physical phantom data. A cylindrical phantom with spheres of diameter 606060~{}60mm, 282828~{}28mm, and 222222~{}22mm with an initial sphere activity concentration of 1.37 kBq/mL was filled at a 10:1 source to background ratio. 34 SPECT acquisitions of the phantom were taken in sequence on a Symbia T2 SPECT/CT system (Siemens Healthineers, USA) with the following settings: $128\times 128$ pixels at $4.8~{}\mathrm{mm}\times 4.8~{}\mathrm{mm}$ resolution, 96 projection angles, high energy collimators, and 60 s acquisition time per projection; although 96 projection angles were acquired, only 32 were used since this was similar to the corresponding 225Ac patient acquisition in the next section. The energy windows selected for TEW additive correction are shown in Table 1. A blank scan with identical parameters was acquired to obtain the mean stray radiation noise in each energy window. Images were reconstructed with MLEM for up to 200 iterations. Two noise levels of the data were considered:
@@ 2.3.2 Physical Phantom Studies
    One of the 34 scans was used for data reconstruction. The number of acquired counts in this scenario is approximately representative of a clinical 225Ac-PSMA-617 scan, where the patient is injected with 888~{}8MBq of activity [23] and scanned for 2.5 min per projection at some time point between 0 hr and 72 hr post injection.
@@
    The counts from all 34 scans were summed together and used for image reconstruction. Because the detectors had an approximately equal radial path around the phantom for each scan, this corresponds to a single scan with a high count rate.
Spherical masks of each hot sphere were drawn using bright edges on the CT image (visible in Figure 5) so that the counts could be computed for each iteration number. Three additional spherical masks of diameter 50 mm were drawn in the background region in order to compute contrast.
@@ 2.3.3 Patient Study
A patient receiving 225Ac-PSMA-617 therapy with an injected activity of 8 MBq was imaged 20.5 hours post injection (first cycle). SPECT data were acquired on a Discovery 670 Pro SPECT/CT (GE Healthcare, USA) with a high energy general purpose collimator. 30 projections (15 per head) were acquired for 150 s with energy windows given by Table 1. Since this GE scanner has a different CDR than the Siemens scanner used in prior examples, PSF data were regenerated in SIMIND using a high energy general purpose GE collimator and another 1DR-MC model was fit as before. In GE scanners, the hexagonal bores in the collimator are rotated by $90^{\circ}$ compared to the Siemens collimator, so the angles in $\Theta_{t}$ were adjusted to compensate for this. The acquired data were reconstructed using 2DG, 2D-MC, and 1DR-MC PSF modeling and MLEM for up to 200 iterations. Three lesion ROIs were segmented by a physician on a pre-therapeutical PET image using the PET Edge+ tool of MIM v7.2.1 (MIM Software Inc., USA). A spherical ROI of diameter 101010~{}10cm was placed in a low uptake region in the center of the patient so that the contrast, defined as the mean uptake ratio between the lesion ROIs and background ROI, could be obtained. The mean number of counts and contrast for each lesion ROI were then evaluated for each iteration of MLEM; in this case, before the statistics were computed for each iteration, a 333~{}3cm FWHM Gaussian filter was applied to the image.
@@ 3 Results
Figure 2 compares the Monte Carlo simulated 2D-MC PSF profiles to the 1DR-MC fit obtained via Equation 8 for 177Lu, 131I, and 225Ac. The 1D-R model reasonable approximates all PSFs, but (i) does not always accurately capture the septal scatter component as seen in the central vertical profiles and (ii) does not capture some additional faint septal penetration tails of 177Lu; this is due to a limitation of the model for only including three tails at specific angles.
@@
Figure 2: Comparison between Monte Carlo simulated PSF data and the corresponding fit from the 1DR-MC model at different source-detector distances. From top to bottom, (i) 2D profiles of Monte Carlo PSF data, (ii) 2D profiles of 1DR-MC model, (iii) central vertical profile of Monte Carlo and fitted PSFs, (iv) central horizontal profile of Monte Carlo and fitted PSFs; the Monte Carlo data is shown in black scatter points, while the fitted data is shown as a solid red line.
@@
Timing benchmarks for PSF modeling using the 1DR-MC, 1DR-MC (FFT), 2D-MC, and 2D-MC (FFT) are shown in Figure 3. CPU implementation yields no benefits with the proposed 1DR-MC model, and performs fastest with FFT-based 2D convolutions and slowest using regular 2D convolutions. The GPU implementation is faster than the CPU implementation for all methods, and yields computational benefits when the proposed 1DR-MC model is used. With a matrix size of $128^{3}$, use of the 1DR-MC method is over three times faster than 2D-MC (FFT) PSF modeling. For small matrix sizes ($64^{3}$) there is no computational speed-up using the 1DR-MC method, and for large matrix sizes ($256^{3}$) the relative time difference between the 1DR-MC and 2D-MC (FFT) begins to decrease, approaching only a two times speed advantage.
@@
Figure 3: CPU and GPU required run time of the 1DR-MC and 2D PSF models for different matrix sizes. Each implementation is evaluated using standard and FFT based convolution. The time shown on top of each bar in units of ms.
Reconstructed MC simulations of the 177Lu, 131I, and 225Ac spherical phantoms are shown in Figure 4, while corresponding RC and contrast statistics (for the last iteration of reconstruction) are shown in Table 2. The 1DR-MC and 2D-MC models yield similar RCs in all cases, and are larger than those obtained using the 2DG model.
Figure 4: Reconstructions of the MC simulations (left) and corresponding RCs for each of the three spheres at each iteration (right). The displayed images correspond to reconstructions after 40 iterations (177Lu), 10 iterations (131I) and 100 iterations (225Ac)
Table 2: Recovery coefficients (or counts) and contrast for various regions of interest in all reconstruction examples using (i) traditional 2D Gaussian PSF modeling (2DG), (ii) 2D kernels (2D-MC), and (iii) the proposed 1DR-MC model that aims to to be consistent with the 2D-MC model while reducing reconstruction times. The low count 225Ac physical phantom corresponds to the first of the 34 acquisitions.
  • 177Lu MC Phantom
    Contrast RC [%]
    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
    2DG 2.68 2.00 1.73 28.4 21.2 18.3
    2D-MC 7.85 5.72 4.23 95.9 69.9 51.7
    1DR-MC 8.19 6.01 4.40 94.6 69.4 50.8
    131I MC Phantom
    Contrast RC [%]
    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
    2DG 8.08 6.94 6.22 75.4 64.7 58.0
    2D-MC 10.41 9.60 8.74 94.1 86.7 79.0
    1DR-MC 10.31 9.49 8.64 94.0 86.5 78.8
    225Ac MC Phantom
    Contrast RC [%]
    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
    2DG 4.55 3.73 3.63 45.5 37.3 36.2
    2D-MC 9.11 8.42 7.05 98.8 91.2 76.5
    1DR-MC 9.03 8.36 6.99 98.7 91.4 76.5
    225Ac Physical Phantom (Low Count)
    Contrast Counts
    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
    2DG 4.01 3.68 3.32 0.0342 0.0314 0.0283
    2D-MC 4.85 4.80 4.36 0.0640 0.0634 0.0575
    1DR-MC 4.88 4.80 4.36 0.0643 0.0632 0.0574
    225Ac Physical Phantom (High Count)
    Contrast Counts
    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
    2DG 4.68 4.31 3.70 1.24 1.14 0.98
    2D-MC 8.15 7.83 5.98 2.52 2.42 1.85
    1DR-MC 8.05 7.69 5.92 2.56 2.45 1.88
    225Ac Patient
    Contrast Counts
    Model Les 1 Les 2 Les 3 Les 1 Les 2 Les 3
    2DG 2.32 2.28 2.80 0.0405 0.0267 0.0660
    2D-MC 3.72 3.50 4.86 0.0405 0.0273 0.0653
    1DR-MC 3.74 3.50 4.92 0.0246 0.0254 0.0355
@@
Reconstructions of physical 225Ac phantom images are shown in Figure 5, and quantitative statistics for (i) the first low count acquisition and (ii) the high count acquisition after 200 iterations are shown in Table 2. Images generated from the 1DR-MC and 2D-MC methods are almost indistinguishable qualitatively. The quality metrics in Table 2 demonstrate similar counts and contrast for the 2D-MC and 1DR-MC models, and significantly lower counts and contrast for the 2DG model. The time required for 200 iterations of MLEM for the low count data was 777777~{}77s (2DG PSF), 299299299~{}299s (1DR-MC PSF) and 748748748~{}748s (2D-MC PSF). All 34 low count acquisitions were also reconstructed and the 2DG, 2D-MC, and 1DR-MC methods were quantitatively compared; the differences in total counts between the 2D-MC and 1DR-MC were $(-0.01\pm 1.20)\%$, $(0.67\pm 0.23)\%$, and $(-0.75\pm 1.25)\%$ for smallest to largest sphere respectively. As a demonstration of how minor these differences are, the variability of counts between the 34 acquisitions for the 2D-MC model were $61.3\%$, $38.5\%$, and $11.1\%$ from smallest to largest sphere, respectively.
@@
Figure 5: Reconstruction of physical 225Ac phantom using various PSF models. Reconstructed SPECT and CT images are superimposed; the boundaries of each sphere can be seen on the CT. The displayed images correspond to reconstruction after 100 iterations; the low count images were additionally post-smoothed using a 3D Gaussian function with a 3 cm FWHM. The plots on the right show the average counts per voxel in the three spheres.
PyTomography reconstructed 225Ac-PSMA-617 patient images are shown in Figure 6, along with a reference 68Ga-PSMA-617 PET scan acquired on a GE Discovery IQ scanner and reconstructed using Q.Clear ($\beta=450$). The time required for 200 iterations of MLEM was $75.1$ s (2DG), 272.0272.0272.0~{}272.0s (1DR-MC) and 704.0704.0704.0~{}704.0s (2D-MC). Qualitatively, images reconstructed using the Gaussian PSF can be observed to have less counts in the lesion ROIs. Quantitative results are shown as line plots in Figure 6, and quality metrics for iteration 200 are shown in Table 2. These results demonstrate (i) indistinguishable images and nearly identical image metrics between the 1DR-MC and 2D-MC models and (ii) higher contrast and count uptake for 2D-MC and 1DR-MC models relative to the 2DG model for all three lesions.
@@
Figure 6: Coronal slice of a 68Ga-PSMA-617 PET image (top row, left) and associated 225Ac-PSMA-617 SPECT data (top row, right), counts in bone lesion ROIs as a function of iteration number (middle row) and signal to background ratios in the bone lesion ROIs (bottom row). 225Ac images were reconstructed in PyTomography using the 2DG, 2D-MC, and 1DR-MC PSF models. For computation of the data in the line plots, the image at each iteration was post-filtered using a 3 cm FWHM 3D Gaussian filter before the statistic was computed. Arrows pointing to the three bone lesions are shown in red (lesion 2 is directly on top of its annotation). There was no uptake in the right kidney due to organ failure.